# Performance-lossless Black-box Model Watermarking

### v1

@@ 3.3.3. Proof
4. Performance-lossless and Secure Watermarking for Language Model
@@
In this section, we provide a concrete example of the branch backdoor-based watermarking protocol for model, using a language model as a case study. Then we further analyze the possible attacks when the threat model faces the watermark protocol during the forensic phase, and propose a more secure instantiation strategy.
4.1. Performance-lossless Watermarking for Language Model
In this section, we take the language model as an example to describe a performance-lossless BBWP based on the unforgeable signature scheme. For the cryptography primitive, we use the ECDSA digital signature algorithm and SHA256 algorithm.
4.1.1. Language Model
Algorithm 1 Backdoor-Activate Function
@@ 4.1.1. Language Model
$x_{tr}\leftarrow x||t$
Output $x_{tr}$
4.1.2. Watermarking Instance
Algorithm 1 defines the Backdoor-Activate function, which is responsible for generating triggers using a digital signature scheme. It combines the message encode function Mac and an embedding function Embed. The input is an arbitrary-length prompt $x$ and the signing key $sk$. It first applies the Mac function to obtain a signature $\sigma$ of the prompt $x$ in token form using the signing key $sk$. Then, it uses the Embed function to convert the signature $\sigma$ (i.e., trigger) into token form, which is concatenated to the end of the original prompt $x$. The final output is a trigger version $x_{tr}$ for $x$.
Unlike image pixels, both text and tokens lack the ideal redundancy to carry triggers. Here we implement a simple Embed function that converts the trigger directly into the same form as the signed message and concatenates them. This approach is feasible without considering the possibility of an attacker detecting and intercepting anomalous queries.
@@ 4.1.2. Watermarking Instance
Algorithms 2 and 3 define the Backdoor-Colloction function. It includes the trigger detection algorithm Detect and the evidence generation algorithm Prove. The Detect function uses the verify algorithm Veri corresponding to the message encode function Mac used in the Backdoor-Activate function. The symmetric key $k$ is the implicit input of the Veri function. It takes a prompt $x_{*}$ as input, converts $x_{*}$ into a binary sequence, and extracts a fixed-length sequence from the end as a potential trigger (i.e., signature), $\sigma$. The remaining part of $x_{*}$ is $x$. Then, the Veri function is used to check whether $\sigma$ is a valid message tag of $x$ under the symmetric key $k$. Determine whether to activate the Prove function based on the trigger detection results. Here we implement a simple Prove function that replaces the model’s inference results directly with statements such as “I am model B from owner A!” The content format can vary, including token sequences, sharp probability distributions, or logical values, depending on the API’s output format.
@@
4.2. Interference Attacks
@@
In practical use, the branch backdoor-based watermarking protocol for model may face targeted interference at tacks against the forensic (i.e., evidence collection) process from attackers when dealing with two types of dishonest behaviors in the threat model. In this section, we discuss the following two attack scenarios:
@@
Scenario 1: A low-cost model B pretends to be a high-cost model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output side. A straightforward way is to reject triggers as illegal inputs. However, the number of triggers is enormous and rejecting a single trigger is reasonable, but the act of rejecting all triggers may itself be suspicious. Therefore, attackers need to take potential trigger requests seriously. When attackers can access high-cost model A, the trigger is detected and then sent to A, returning the corresponding copyright evidence, thus successfully deceiving at a critical moment in the forensic process. When attackers cannot access A, they detect triggers and attempt to forge evidence.
@@
Scenario 2: Lie about having a self-developed high-cost model B but it is actually model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output end. When attackers have a low-cost alternative model C, they send the detected triggers to it to return corresponding outputs, indicating that the triggers have no effect on B, thus demonstrating that there is no copyright dispute. When attackers do not have model C available, they send the detected triggers to A to obtain corresponding copyright evidence and perform an evidence erasure attack to achieve the same purpose. Additionally, attackers can also detect copyright evidence at the output side and return the same alternative or erased output.
4.3. Analysis for Interference Attacks
@@
  1. (1)
    Triggers and copyright evidence are stealthy and diverse, making it difficult for attackers to detect and intercept;
  2. (2)
@@
Notably, an attacker can distort the trigger to disable it, thereby bypassing the evidence collection process. However, in the text generation task of language models, tampering with the input brings more uncertainty than erasing the known output, which may seriously affect the attacker’s application service quality. Therefore, Principle (3) can be relaxed to only emphasize the robustness of copyright evidence.
Obviously, the current output of the Backdoor-Activate function, algorithm 1, lacks stealth. The Embed function directly converts the signature sequence into a token sequence with no natural semantics, resulting in garbled characters after being appended to the original prompt. It can be easily detected using simple perplexity detection. The inference output content replaced by the Prove function, algorithm 3, is also easy to detect and forged by attackers, lacking stealth and robustness.
4.4. Improving Security of Watermarking
In this section, we improve the previously proposed simple structure to resist potential interference attacks during evidence collection.
4.4.1. Stealthy Input and Output
In this section, we show how to improve the stealth and robustness of the Backdoor-Activate, Detect, and Prove functions constructed in section 4.1.2 by integrating existing no-box watermark schemes, including a robust multi-bit watermark method (Yoo et al., 2023) and a simple multi-bit watermark method modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023).
Algorithm 4 defines the stealthy version of the Embed function in the Backdoor-Activate function. A stealthy Embed function is implemented using the watermarked text generation algorithm of the semantic-preserving no-box watermark scheme modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023). It takes an arbitrary-length prompt $x$ and the signature $\sigma$ of $x$ as input, with the embedding key $ek_{in}$ as an implicit input. The binary sequence of the signature $\sigma$ is embedded into a newly generated text that has $x$ as context (i.e., initial history). Then, the newly generated text is appended to $x$ to produce $x_{tr}$.
Algorithm 4 Activate-Stealthy Embed Function
Prompt $x$, Signature $\sigma$, Embedding Key $ek$, Model $\mathcal{M}$, Vocabulary $\mathcal{V}$
Triggered Prompt $x_{tr}$
@@ 4.4.1. Stealthy Input and Output
$x_{tr}\leftarrow\text{tokenizer.decode}(history)$
Output $x_{tr}$
Algorithm 5 Detect-Stealthy Extract Function
Prompt $x_{*}$, Verifying Public Key $pk$, Vocabulary $\mathcal{V}$
@@
Algorithm 5 redefines the Extract function in the Detect function to adapt to the improved stealthy Embed function in the Backdoor-Activate function. The extraction operation is implemented using the watermarked text extraction algorithm of the semantic-preserving no-box watermark scheme modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023). It takes the prompt $x_{*}$ as input, with the embedding key $ek$ as an implicit input. The prompt $x_{*}$ is hashed token-by-token to extract the potential signature $\sigma$ and the original prompt $x$.
Algorithm 6 defines the stealthy and robust version of the Prove function. A stealthy and robust Prove function is implemented using the logit modification algorithm of the semantic-preserving and robust no-box watermark scheme (Yoo et al., 2023). It takes the copyright information $c$ and the original model’s inference result $l$ as input, with the embedding key $ek_{out}$ as an implicit input. Then, embed $c$ into $l$ to obtain a modified inference result $l_{c}$, which is then output as the final inference result.
Algorithm 6 Prove-Stealthy Embed Function
Trigger Detection Result $r$, Original Logit $y$, Prefix $prefix$, Copyright Information $c$, Embedding State $(\delta,j)$, Embedding Key $ek$, Vocabulary $\mathcal{V}$
@@
     $seed\leftarrow\text{Hash1}(prefix)$
@@
     $start\leftarrow\text{Hash2}(prefix)\mod j$
@@
4.4.2. Security Analysis
The triggers and copyright evidence are stealthy and diverse. The semantic-preserving watermarking scheme ensures that the triggers and copyright evidence are stealthy. The prompt space provides a variety of options for the original prompt $x$. The signing function can be a probabilistic algorithm that provides multiple signatures for a message. When $x$ has high entropy, it can generate diverse triggers with consistent context but varied expressions. Diverse triggers can provide various copyright evidence. Therefore, the triggers and copyright evidence are diverse.
@@ 4.4.2. Security Analysis
Copyright evidence is robust. Taking advantage of the inherent robustness of the stealthy watermarking scheme, a certain degree of insertions, deletions, and replacements can be tolerated. It is worth noting that if an attacker attempts to make large-scale and high-quality modifications to copyright evidence, the costs will outweigh the benefits.
5. Discussion and Limitations
5.1. Discussion
5.1.1. Replacement Attacks
@@
The construction in this paper is based on message authentication technology in cryptography, which naturally raises the question of replay attacks, a threat faced by MAC. In this paper, due to the vulnerability of MAC to replay attacks, attackers could execute trigger replay attacks. However, this type of attack is of limited practical significance.
@@
We envision an alternative replay attack called the “copyright evidence replay attack” for the proposed scheme. While attackers cannot forge valid copyright evidence for a new trigger, they can search for semantically similar triggers they have encountered before and reuse their copyright evidence as the response to the new trigger. This attack highlights a weakness of our scheme, that is, the correlation between trigger input and evidence output is relatively weak, and semantic relevance alone is insufficient.
@@
To solve this problem, we draw inspiration from solutions in MAC technology to incorporate timestamps into copyright evidence to maintain temporal correlation beyond semantic correlation. Another idea is to hash the carried tag $\sigma$ and use it as the embedded key $ek$ for the Prove function. This strong connection between triggered prompt and copyright evidence increases resistance to forgery and replay attacks.
5.1.2. Generalization
@@
This paper introduces a general watermarking protocol framework for model in black-box scenarios. We give a branch indicator based on the MAC scheme. In fact, the MAC scheme can be replaced by any cryptographic primitives with unforgeability security. In other words, the underlying cryptographic primitives are flexible and extensible. Besides, while the watermarking example is constructed for pre-trained language models in text generation tasks, the framework is applicable to various media formats, such as images and audio, and is not limited by network structures or tasks. As long as the actual application requirements are met, appropriate trigger embedding and extraction functions and representations of copyright evidence can be designed.
5.2. Limitations
@@
The performance-lossless property defined in this paper for model watermarking is contradictory to the defense against model extraction attacks in black-box scenarios. In a model extraction attack, attackers aim to obtain a training dataset by querying the victim model to train an alternative model capable of stealing the functionality of the original model. Attackers who have difficulty obtaining triggers have no opportunity to access the model’s inference results modified by the no-box scheme under the forensic state. Defense against model extraction attacks often relies on having a learnable watermark pattern within the (query, response) data available to attackers. When an attacker successfully performs a model extraction attack on a black-box model, it means that the attacker can easily trigger the API’s forensic state, which means that the target model is far from the performance-lossless property.
@@
However, model extraction attacks are indeed a significant threat to protecting model intellectual property in a black-box scenario. In practice, we can refer to the approach used in the work (Szyller et al., 2021), where watermark injection operations are randomly applied to queries under the service state, sacrificing a certain degree of model performance to meet security requirements.
6. Conclusion

### v2

@@ III-E3 Hash vs MAC vs Signature
Therefore, we chose an efficient and secure MAC to construct a performance-lossless branch watermarking scheme for model.
IV Performance-lossless and Secure Watermarking for Language Model
@@
In this section, we provide a concrete example of the branch backdoor-based watermarking protocol for model using a language model as a case study. Then, we further analyze the possible attacks when the threat model faces the watermarking protocol during the forensic phase and propose a more secure instantiation strategy.
Algorithm 1 TriggerGen Function
@@
$x_{tri}\leftarrow x||t$
Output $x_{tri}$
@@
IV-A Performance-lossless Watermarking for Language Generation Model
In this section, we take the language model as an example to describe a performance-lossless BranchWM based on the unforgeable MAC scheme.
IV-A1 Language Model
IV-A2 Watermarking Instance
Algorithm 1 defines the TriggerGen function, which is responsible for generating triggers using a digital signature scheme. It combines the message encoding function Mac and an embedding function Embed. The input is an arbitrary-length prompt $x$ and the signing key $sk$. It first applies the Mac function to obtain a signature $\sigma$ of the prompt $x$ in token form using the signing key $sk$. Then, it uses the Embed function to convert the signature $\sigma$ (i.e., trigger) into token form, which is concatenated to the end of the original prompt $x$. The final output is a trigger version $x_{tri}$ for $x$.
Unlike image pixels, both text and tokens lack the ideal redundancy to carry triggers. Here, we implement a simple Embed function that converts the trigger directly into the same form as the signed message and concatenates them. This approach is feasible without considering the possibility of an attacker detecting and intercepting anomalous queries.
The Verify algorithm is used to confirm whether the response of the suspicious API to the trigger query is the statement set in the Prove function.
Algorithms 2 and 3 define the Branch function. It includes the trigger detection algorithm Detect and the evidence generation algorithm Prove. The Detect function uses the verify algorithm Veri corresponding to the message encoding function Mac used in the TriggerGen function. The symmetric key $k$ is the implicit input of the Veri function. It takes a prompt $x_{*}$ as input, converts $x_{*}$ into a binary sequence, and extracts a fixed-length sequence from the end as a potential trigger (i.e., signature), $\sigma$. The remaining part of $x_{*}$ is $x$. Then, the Veri function is used to check whether $\sigma$ is a valid message tag of $x$ under the symmetric key $k$. Determine whether to activate the Prove function based on the trigger detection results. Here we implement a simple Prove function that replaces the model’s inference results directly with statements such as “I am model B from owner A!” The content format can vary, including token sequences, sharp probability distributions, or logical values, depending on the API’s output format.
Figure 4: Interference attacks in two scenarios: (1) The attacker uses a low-cost model B to feign a high-cost model A and tries to filter the input or forge the output; (2) The attacker lies about having self-developed high-cost model B which is actually model A and tries to filter the input or erase the watermark signal in the output.
IV-B Interference Attacks
@@
In practical use, the branch backdoor-based watermarking protocol for model may face targeted interference at tacks against the forensic (i.e., evidence collection) process from attackers when dealing with two types of dishonest behaviors in the threat model. Figure 4 shows the following two attack scenarios:
@@
Scenario 1: A low-cost model B pretends to be a high-cost model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output side. A straightforward way is to reject triggers as illegal inputs. However, the number of triggers is enormous and rejecting a single trigger is reasonable, but the act of rejecting all triggers may itself be suspicious. Therefore, attackers need to take potential trigger requests seriously. When attackers can access high-cost model A, the trigger is detected and then sent to A, returning the specific copyright evidence, thus successfully deceiving at a critical moment in the forensic process. When attackers cannot access A, they detect triggers and attempt to forge evidence.
@@
Scenario 2: Lie about having self-developed high-cost model B, but it is actually model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output end. When attackers have a low-cost alternative model C, they send the detected triggers to it to return specific outputs, indicating that the triggers have no effect on B, thus demonstrating that there is no copyright dispute. When attackers do not have model C available, they send the detected triggers to A to obtain specific copyright evidence and perform an evidence erasure attack to achieve the same purpose. Additionally, attackers can also detect copyright evidence at the output side and return the same alternative or erased output.
IV-C Analysis for Interference Attacks
@@
  1. 1.
    Triggers and copyright evidence are concealed and diverse, making it difficult for attackers to detect and intercept;
  2. 2.
@@
Notably, an attacker can distort the trigger to disable it, thereby bypassing the evidence collection process. However, in the text generation task of language models, tampering with the input brings more uncertainty than erasing the known output, which may seriously affect the attacker’s original service quality. Therefore, Principle (3) can be relaxed to only emphasize the robustness of copyright evidence.
Obviously, the current output of the TriggerGen function, algorithm 1, lacks stealth. The Embed function directly converts the signature sequence into a token sequence with no natural semantics, resulting in garbled characters after being appended to the original prompt. It can be easily detected using simple perplexity detection. The inference output content replaced by the Prove function, algorithm 3, is also easy to detect and forged by attackers, lacking stealth and robustness.
IV-D Improving Security of Watermarking
In this section, we improve upon the previously proposed simple structure to resist potential interference attacks during evidence collection.
We show how to improve the concealment and robustness of the TriggerGen, Detect, and Prove functions constructed in section IV-A2 by integrating existing no-box watermarking schemes, including a robust multi-bit watermarking method [50] and a simple multi-bit watermarking method modified from Kirchenbauer et al.’s work [45].
Algorithm 4 Concealed TriggerGen: Embedding Function
Prompt $x$, Signature $\sigma$, Model $\mathcal{M}$, Vocabulary $\mathcal{V}$
Triggered Prompt $x_{tri}$
@@ IV-D Improving Security of Watermarking
$x_{tri}\leftarrow\text{tokenizer.decode}(history)$
Output $x_{tri}$
Algorithm 4 defines the concealed version of the Embed function in the TriggerGen function. A concealed Embed function is implemented using the watermarked text generation algorithm of the semantic-preserving no-box watermarking scheme modified from Kirchenbauer et al.’s work [45]. It takes an arbitrary-length prompt $x$ and the signature $\sigma$ of $x$ as input, with the embedding key $ek_{in}$ as an implicit input. The binary sequence of the signature $\sigma$ is embedded into a newly generated text that has $x$ as context (i.e., initial history). Then, the newly generated text is appended to $x$ to produce $x_{tri}$. Table II displays the simple triggers from algorithm 1 and the concealed triggers generated from algorithm 4. It shows that the latter maintains semantic readability and has good concealment, while the former is almost garbled.
Algorithm 5 Concealed Detect: Extract Function
Prompt $x_{*}$, Vocabulary $\mathcal{V}$
@@
Algorithm 6 Concealed Prove: Embedding Function
Trigger Detection Result $r$, Original Logit $y$, Signature $\sigma$, Prefix $prefix$, Copyright Information $c$, Embedding State $(\delta,j)$, MAC Key $ek$, Vocabulary $\mathcal{V}$
@@
$\sigma^{\prime}\leftarrow\text{Mac}(k,\sigma)$
$\sigma^{\prime}_{1},\sigma^{\prime}_{2}\leftarrow\sigma^{\prime}$triangleright Divide $\sigma^{\prime}$ into two parts from the middle
@@
     $seed\leftarrow\text{Hash1}(\sigma^{\prime}_{1}||prefix)$
@@
     $start\leftarrow\text{Hash2}(\sigma^{\prime}_{2}||prefix)\mod j$
@@
Algorithm 5 redefines the Extract function in the Detect function to adapt to the improved concealed Embed function in the TriggerGen function. The extraction operation is implemented using the watermarked text extraction algorithm of the semantic-preserving no-box watermarking scheme modified from Kirchenbauer et al.’s work [45]. It takes the prompt $x_{*}$ as input, with the embedding key $ek$ as an implicit input. The prompt $x_{*}$ is hashed token-by-token to extract the potential signature $\sigma$ and the original prompt $x$.
Algorithm 6 defines the concealed and robust version of the Prove function. A concealed and robust Prove function is implemented using the logit modification algorithm of the semantic-preserving and robust no-box watermarking scheme [50]. The pseudocode of the algorithm shows the modification of the original model inference result $y_{i}$ by the Prove function during the single-step token generation process. At each step of token generation, the copyright information $c$ is partially embedded into the modified reasoning result $y_{i}^{\prime}$. This algorithm is called repeatedly until the complete response content is generated. The complete response process takes the copyright information $c$ and the original model’s inference result $y$ as input, with the embedding key $ek_{out}$ as an implicit input. Then, $c$ is embedded into $y$ to obtain a modified inference result $y_{c}$, which is then output as the final inference result.
The complete Verify algorithm includes copyright information extraction and verification. The extraction function can be implemented using the extraction algorithm [50] corresponding to Algorithm 6. During the forensics process, infringement is confirmed when the Verify algorithm successfully extracts and verifies the suspect API’s response to the trigger.
IV-D1 Security Analysis
The triggers and copyright evidence are concealed and diverse. The semantic-preserving watermarking scheme ensures that the triggers and copyright evidence are concealed. The prompt space provides a variety of options for the original prompt $x$. The signing function can be a probabilistic algorithm that provides multiple signatures for a message. When $x$ has high entropy, it can generate diverse triggers with consistent context but varied expressions. Diverse triggers can provide various copyright evidence. Therefore, the triggers and copyright evidence are diverse.
@@ IV-D1 Security Analysis
Copyright evidence is robust. Taking advantage of the inherent robustness of the concealed watermarking scheme, a certain degree of insertions, deletions, and replacements can be tolerated. Notably, if an attacker attempts to make large-scale and high-quality modifications to copyright evidence, the costs will outweigh the benefits.
Table II: Examples of simple triggers and concealed triggers (truncated for display). Take the single sentence in the story reasoning dataset ROCStories [56] as pormpt, HMAC_SHA512 generates the corresponding trigger pattern, and uses the tokenizer and model of OPT-1.3B [57] to generate triggers.
Prompt Simple Trigger Concealed Trigger
[…]
V Discussion and Limitations
V-A Discussion
V-A1 Replay Attacks
@@
The construction in this paper is based on message authentication technology in cryptography, which naturally raises the question of replay attacks, a threat faced by MAC. In this paper, due to the vulnerability of MAC to replay attacks, attackers can execute trigger replay attacks. However, this type of attack is of limited practical significance.
@@
We envision an alternative replay attack called the “copyright evidence replay attack” for the proposed scheme. While attackers cannot forge valid copyright evidence for a new trigger, they can search for semantically similar triggers they have encountered before and reuse their copyright evidence as a response to the new trigger. This attack highlights a weakness of the BranchWM; that is, the correlation between the trigger input and the evidence output is relatively weak, and semantic relevance alone is insufficient.
@@
To solve this problem, we draw inspiration from MAC technology solutions to incorporate timestamps into copyright evidence to maintain temporal correlation beyond semantic correlation. Another idea is to hash the carried tag $\sigma$ and use it as the embedded key $ek$ for the Prove function, as shown in Algorithm 6. This strong connection between triggered prompt and copyright evidence increases resistance to forgery and replay attacks.
V-A2 Generalization
@@
This paper introduces a general watermarking protocol framework for model in black-box scenarios. We give a branch indicator based on the MAC scheme. In fact, the MAC scheme can be replaced by any cryptographic primitive with unforgeability security to adapt to specific scenarios. In other words, the underlying cryptographic primitives are flexible and extensible. In addition, while the watermarking example is constructed for pre-trained language models in text generation tasks, the framework is applicable to various media formats, such as images and audio, and is not limited by network structures or tasks. As long as the actual application requirements are met, appropriate trigger embedding and extraction functions and representations of copyright evidence can be designed.
Image Classification Task. Different from the text generation task exemplified in Section 4, images are more redundant than text, and their embeddable space can meet the length requirements of secure cryptographic primitives. The following is a simple example of performance-lossless branch watermarking for an image classification model: within the TriggerGen algorithm, the specific pixel plane of the original image is input into the Mac algorithm to generate a trigger pattern, and then a simple Least Significant Bit ( LSB ) replacement method is used to embed the pattern into the image to obtain a trigger image; within the Detect algorithm, the tag of the specific pixel plane is recalculated and then compared with the LSB plane. If they are the same, the image is a trigger image; within the Prove algorithm, the model output (i.e., prediction label) is replaced with the last bit value of the trigger pattern; the Verify algorithm detects whether the label of the trigger image predicted by the suspicious API is equal to the last bit value it carries.
V-B Limitations
V-B1 Model Extraction Attacks
@@
The performance-lossless property defined in this paper for model watermarking contradicts the defense against model extraction attacks in black-box scenarios. In a model extraction attack, attackers aim to obtain a training dataset by querying the victim model to train an alternative model capable of stealing the functionality of the original model. Attackers who have difficulty obtaining triggers have no opportunity to access the model’s inference results modified by the no-box scheme under the forensic state. Defense against model extraction attacks often relies on having a learnable watermark pattern within the (query, response) data available to attackers. When an attacker successfully performs a model extraction attack on a black-box model, the attacker can easily trigger the API’s forensic state, which means that the target model is far from being performance-lossless.
@@
However, model extraction attacks are indeed a significant threat to protecting model intellectual property in a black-box scenario. In practice, we can refer to Szylle et al.’s work [58], where watermark injection operations are randomly applied to queries under the service state, sacrificing a certain degree of model performance to meet security requirements.
V-B2 Implementation
Length of trigger. Due to the length limitation of the trigger pattern based on a secure cryptographic primitive and the small capacity of existing text watermarking, the examples given in this paper are only suitable for text generation tasks that support long text input. How to implement performance-lossless branch watermarking for short text input language models is a direction that requires further practical design considerations.
Detection of machine-generated text. In addition, the trigger text generated by the proposed instance is essentially machine-generated text. Although the machine-generated trigger text can maintain good semantic readability, it is not enough to achieve excellent concealment. Concealment also involves the rationality of query behavior. The query received by the generated model is usually natural text. Obviously, the machine-generated trigger text is easily detected by the existing excellent machine text detection algorithms. However, it is extremely difficult for the detector to reach 0 false alarm. A model that provides services always faces a large number of queries. When the attacker uses the machine text detector to filter the forensic attempt of the model owner, it will also lose a large number of normal user queries, which will seriously affect the piracy service. Therefore, the cost of this kind of forensic interference attack is extremely high. Of course, we still look forward to natural text hiding methods with large capacity and efficient extraction in the future to expand upon the current trigger generation methods.
VI Conclusion